Jesse works in a bakery. He can make 12 pics in 96 minutes. In the space below, write an equation that can be used to find how many pics(p) Jesse would be able to make in (t) minutes.
Equation
If Jesse makes pics for 8 hours, how many pics could he make?
show your work
Thanks
7th grader in NY

Answer by stanbon(48558)

Jesse works in a bakery. He can make 12 pics in 96 minutes. In the space below, write an equation that can be used to find how many pics(p) Jesse would be able to make in (t) minutes.
----------
12 pics in 96 min means 1 pic every 8 minutes
EQUATION:
1/8 pic/min
f(t)=(1/8)t
-------
f(t) stands for number of pictures produces
t stands for number of minutes
----------------
If Jesse makes pics for 8 hours, how many pics could he make?
8 hrs is 8*60 = 480 minutes.
f(480) = (1/8)480
f(480)= 60
He would produce 60 pictures in 8 hrs.
----------
Cheers,
Stan H.